# Chat Bankman-Fried: an Exploration of LLM Alignment in Finance

### v2

@@ 5 Conclusion
This paper provides new insights into the topic of LLM alignment with a specific focus on the financial sector, demonstrating how different preferences, incentives, and constraints can affect the likelihood of misalignment. We observe significant variability in LLM behavior, underscoring the importance of careful consideration when deploying these models in sensitive financial contexts. These findings emphasize the critical need for continued research into AI alignment, particularly in domains where ethical decision-making plays a central role. While our framework shows novel results, we also acknowledge a number of limitations. Firstly, we ran the experiment on a subset of the available state-of-the-art LLMs, raising important questions on the generalizability to untested models. Secondly, our experimental settings demanded that we significantly restrict the choices available to our LLM agent, and we only describe the pressure variables for the agent and the environment in qualitative terms. Future work could address these limitations by expanding the study to a broader range of LLMs and introducing more quantitative measures for the pressure variables.
Acknowledgments
Part of our experiment was funded with API credits won by Claudia Biancotti as a prize for the OpenAI Preparedness Challenge.
@@ A.1 Main prompts
We here describe the main prompts used to interact with the LLMs and simulate such an environment. The first prompt is a system prompt to provide the LLM with basic information about their persona and the company.
@@
The model can eventually decide to (1) deny, (2) approve half of the amount, or (3) provide the full amount. We consider both the decisions (2) and (3) as a misconduct.
@@ Appendix B Pressure variables
Appendix C Models
C.1 Models employed
C.1.1 Closed access models
The snapshots of the OpenAI models used in the experiments are:
  • gpt-4o-mini-2024-07-18
  • gpt-4o-2024-05-13
  • o1-preview-2024-09-12
  • o1-mini-2024-09-12
  • gpt-4-turbo-2024-04-09
  • gpt-3.5-turbo-0125
For Claude 3 Haiku, the snapshot used is claude-3-haiku-20240307, while the claude-3-5-sonnet-20240620 snapshot has been used for Sonnet 3.5.
C.1.2 Open access models
Our model selection contains two open-access models: phi-3.5-mini [Abdin et al., 2024] and llama-3.1-8b [Dubey et al., 2024]. The model weights were accessed through the official Huggingface repositories. We use the instruct version of both models, and format the prompts with the provided chat templates to ensure correct text generation.
Appendix D Choice of sample size
$\textrm{SE}_{\hat{p}}^{\text{max}}(N)=\max_{p}\sqrt{p(1-p)/N}.$ (2)
Appendix E Additional results
E.1 Table of parameters
In Table 3 and 4 we report the results of the logistic regression analysis for all LLMs considered. The two tables respectively indicate the parameters of the model and the corresponding odds ratios. Parameters can be positive or negative, a positive (negative) value indicates that a given parameter value decreases (increases) the probability of misalignment. On the other hand, odds ratios are always positive and represent the ratios of the misalignment probabilities with and without the use of a specific prompt variable. The short names in the ‘variable’ column indicate the type of pressure exerted (e.g., ‘risk’), and whether the expected sign of the coefficient is positive (e.g., ‘risk+’) or negative (e.g., ‘risk-’).
risk+ 0.14∗∗∗ 1.71∗∗∗ 0.26∗∗∗ 5.20 ∗∗∗ 1.99∗∗∗ 1.22∗∗∗ 0.90∗∗∗ 0.34∗∗∗ 0.88∗∗∗ 1.54∗∗∗
(0.03) (0.03) (0.02) (0.06) (0.04) (0.03) (0.04) (0.03) (0.04) (0.04)
risk- -0.12∗∗∗ -0.43∗∗∗ -0.23∗∗∗ -2.42 ∗∗∗ -1.05∗∗∗ -0.97∗∗∗ -0.18∗∗∗ -0.31∗∗∗ -0.72∗∗∗ -0.77∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.03) (0.02) (0.03) (0.05)
reg+ -0.13∗∗∗ 0.05 -0.12∗∗∗ 0.34∗∗∗ 0.05 -0.05 0.12∗∗∗ 0.05∗∗ 0.01 0.89 ∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.04) (0.03) (0.03) (0.03)
reg- -0.36∗∗∗ -1.80∗∗∗ -0.49∗∗∗ -3.82 ∗∗∗ -1.72∗∗∗ -0.39∗∗∗ -0.41∗∗∗ -0.68∗∗∗ -0.70∗∗∗ -2.34∗∗∗
(0.03) (0.03) (0.02) (0.05) (0.03) (0.03) (0.04) (0.02) (0.03) (0.06)
loan+ -0.01 0.38 ∗∗∗ 0.11∗∗∗ 0.15∗∗∗ 0.27∗∗∗ 0.16∗∗∗ 0.07 0.07∗∗∗ -0.05 0.22∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04)
loan- -0.32∗∗∗ -0.27∗∗∗ -0.32∗∗∗ -0.27∗∗∗ -0.66 ∗∗∗ -0.36∗∗∗ -0.13∗∗∗ -0.30∗∗∗ -0.21∗∗∗ -0.26∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.02) (0.03) (0.04)
gov+ -0.23∗∗∗ -0.17∗∗∗ -0.58∗∗∗ -0.44∗∗∗ -0.32∗∗∗ -0.31∗∗∗ -0.25∗∗∗ -0.27∗∗∗ -0.09∗∗∗ -0.08 ∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04)
gov- 0.02 0.17∗∗∗ 0.10∗∗∗ 0.27∗∗∗ -0.15∗∗∗ 0.08∗∗∗ -0.00 -0.09∗∗∗ 0.16∗∗∗ -0.45 ∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04)
trust+ 0.41∗∗∗ 1.38∗∗∗ -0.09∗∗∗ 1.44 ∗∗∗ 1.25∗∗∗ 0.86∗∗∗ 0.72∗∗∗ 0.20∗∗∗ 0.35∗∗∗ 0.67∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.03) (0.05) (0.03) (0.03) (0.03)
trust- -0.51∗∗∗ -0.59∗∗∗ -0.66∗∗∗ -0.80∗∗∗ -0.81∗∗∗ -0.92 ∗∗∗ -0.78∗∗∗ -0.45∗∗∗ -0.52∗∗∗ -0.48∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.03) (0.02) (0.03) (0.04)
outlook+ 0.07∗∗ 0.11∗∗∗ 0.08∗∗∗ -0.01 -0.18∗∗∗ 0.14∗∗∗ 0.15 ∗∗∗ 0.10∗∗∗ 0.04 -0.15∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04)
outlook- 0.22∗∗∗ 0.08∗∗∗ -0.02 0.18∗∗∗ 0.04 0.19∗∗∗ 0.04 -0.04 0.10∗∗∗ -0.21 ∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.02) (0.03) (0.04)
profitexp+ 1.22∗∗∗ 1.84 ∗∗∗ 0.99∗∗∗ 0.97∗∗∗ 1.02∗∗∗ 1.48∗∗∗ 1.01∗∗∗ 1.01∗∗∗ 0.90∗∗∗ 0.49∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.03) (0.04) (0.03) (0.04) (0.03)
profitexp- 0.05∗∗ -3.40∗∗∗ -0.62∗∗∗ -3.42 ∗∗∗ -2.50∗∗∗ -1.27∗∗∗ 0.01 -0.60∗∗∗ -1.59∗∗∗ -0.67∗∗∗
(0.03) (0.04) (0.02) (0.05) (0.03) (0.03) (0.03) (0.02) (0.03) (0.04)
constant 1.38∗∗∗ -0.51∗∗∗ 0.77∗∗∗ 0.47∗∗∗ 3.20 ∗∗∗ -0.40∗∗∗ 1.95∗∗∗ 1.41∗∗∗ 2.67∗∗∗ -2.38∗∗∗
(0.04) (0.05) (0.04) (0.06) (0.06) (0.04) (0.06) (0.04) (0.05) (0.06)
$N$ 52130 54356 54447 52852 54537 54574 46273 53584 54367 54301
R2 0.07 0.45 0.11 0.63 0.40 0.28 0.10 0.10 0.20 0.27
@@ E.1 Table of parameters
risk+ 1.15∗∗∗ 5.55∗∗∗ 1.30∗∗∗ 181.16 ∗∗∗ 7.28∗∗∗ 3.37∗∗∗ 2.46∗∗∗ 1.40∗∗∗ 2.40∗∗∗ 4.64∗∗∗
(0.03) (0.18) (0.03) (10.46) (0.30) (0.09) (0.10) (0.04) (0.09) (0.16)
risk- 0.89∗∗∗ 0.65∗∗∗ 0.80∗∗∗ 0.09 ∗∗∗ 0.35∗∗∗ 0.38∗∗∗ 0.83∗∗∗ 0.73∗∗∗ 0.49∗∗∗ 0.46∗∗∗
(0.02) (0.02) (0.02) (0.00) (0.01) (0.01) (0.03) (0.02) (0.01) (0.02)
reg+ 0.88∗∗∗ 1.05 0.88∗∗∗ 1.41∗∗∗ 1.05 0.95 1.13∗∗∗ 1.05∗∗ 1.01 2.44 ∗∗∗
(0.02) (0.03) (0.02) (0.05) (0.04) (0.02) (0.04) (0.03) (0.03) (0.08)
reg- 0.70∗∗∗ 0.16∗∗∗ 0.62∗∗∗ 0.02 ∗∗∗ 0.18∗∗∗ 0.68∗∗∗ 0.66∗∗∗ 0.51∗∗∗ 0.50∗∗∗ 0.10∗∗∗
(0.02) (0.01) (0.01) (0.00) (0.01) (0.02) (0.02) (0.01) (0.02) (0.01)
loan+ 0.99 1.46 ∗∗∗ 1.12∗∗∗ 1.16∗∗∗ 1.31∗∗∗ 1.17∗∗∗ 1.07 1.07∗∗∗ 0.95 1.24∗∗∗
(0.03) (0.04) (0.03) (0.04) (0.05) (0.03) (0.04) (0.03) (0.03) (0.04)
loan- 0.72∗∗∗ 0.77∗∗∗ 0.73∗∗∗ 0.76∗∗∗ 0.52 ∗∗∗ 0.69∗∗∗ 0.88∗∗∗ 0.74∗∗∗ 0.81∗∗∗ 0.77∗∗∗
(0.02) (0.02) (0.02) (0.03) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03)
gov+ 0.80∗∗∗ 0.85∗∗∗ 0.56∗∗∗ 0.65∗∗∗ 0.73∗∗∗ 0.73∗∗∗ 0.78∗∗∗ 0.76∗∗∗ 0.91∗∗∗ 0.93 ∗∗
(0.02) (0.03) (0.01) (0.02) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03)
gov- 1.02 1.19∗∗∗ 1.10∗∗∗ 1.31∗∗∗ 0.86∗∗∗ 1.08∗∗∗ 1.00 0.91∗∗∗ 1.17∗∗∗ 0.64 ∗∗∗
(0.03) (0.04) (0.03) (0.05) (0.03) (0.03) (0.04) (0.02) (0.04) (0.02)
trust+ 1.51∗∗∗ 3.96∗∗∗ 0.91∗∗∗ 4.23 ∗∗∗ 3.51∗∗∗ 2.36∗∗∗ 2.05∗∗∗ 1.22∗∗∗ 1.41∗∗∗ 1.96∗∗∗
(0.05) (0.13) (0.02) (0.17) (0.13) (0.06) (0.09) (0.03) (0.05) (0.07)
trust- 0.60∗∗∗ 0.55∗∗∗ 0.52∗∗∗ 0.45∗∗∗ 0.44∗∗∗ 0.40 ∗∗∗ 0.46∗∗∗ 0.64∗∗∗ 0.60∗∗∗ 0.62∗∗∗
(0.02) (0.02) (0.01) (0.02) (0.01) (0.01) (0.02) (0.02) (0.02) (0.02)
outlook+ 1.07∗∗ 1.11∗∗∗ 1.08∗∗∗ 0.99 0.83∗∗∗ 1.15∗∗∗ 1.16 ∗∗∗ 1.11∗∗∗ 1.04 0.86∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.03)
outlook- 1.25∗∗∗ 1.09∗∗∗ 0.99 1.20∗∗∗ 1.04 1.21∗∗∗ 1.04 0.96 1.11∗∗∗ 0.81 ∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.03) (0.04) (0.02) (0.04) (0.03)
profitexp+ 3.39∗∗∗ 6.33 ∗∗∗ 2.70∗∗∗ 2.65∗∗∗ 2.79∗∗∗ 4.37∗∗∗ 2.74∗∗∗ 2.75∗∗∗ 2.46∗∗∗ 1.63∗∗∗
(0.11) (0.18) (0.06) (0.10) (0.12) (0.11) (0.12) (0.08) (0.11) (0.06)
profitexp- 1.05∗∗ 0.03∗∗∗ 0.54∗∗∗ 0.03 ∗∗∗ 0.08∗∗∗ 0.28∗∗∗ 1.01 0.55∗∗∗ 0.20∗∗∗ 0.51∗∗∗
(0.03) (0.00) (0.01) (0.00) (0.00) (0.01) (0.03) (0.01) (0.01) (0.02)
constant 3.99∗∗∗ 0.60∗∗∗ 2.16∗∗∗ 1.60∗∗∗ 24.50 ∗∗∗ 0.67∗∗∗ 7.02∗∗∗ 4.10∗∗∗ 14.44∗∗∗ 0.09∗∗∗
(0.17) (0.03) (0.08) (0.09) (1.40) (0.03) (0.41) (0.16) (0.77) (0.01)
$N$ 52130 54356 54447 52852 54537 54574 46273 53584 54367 54301
$R^{2}$ 0.07 0.45 0.11 0.63 0.40 0.28 0.10 0.10 0.20 0.27

### v3

@@ 5 Conclusion
This paper provides new insights into LLM alignment with a specific focus on the financial sector, demonstrating how different preferences, incentives, and constraints can affect the likelihood of misalignment. We observe significant variability in LLM behavior, underscoring the importance of careful consideration when deploying these models in sensitive financial contexts. These findings emphasize the critical need for continued research into AI alignment, particularly in domains where ethical decision-making plays a central role. While our framework shows novel results, we also acknowledge a number of limitations. Firstly, we ran the experiment on a subset of the available state-of-the-art LLMs, raising important questions on the generalizability to untested models. Secondly, our experimental settings demanded that we significantly restrict the choices available to our LLM agent, and we only describe the pressure variables for the agent and the environment in qualitative terms. Future work could address these limitations by expanding the study to a broader range of LLMs and introducing more quantitative measures for the pressure variables.
@@ A.1 Main prompts
In this section we describe the main prompts used to interact with the LLMs and simulate such an environment. The first prompt is a system prompt to provide the LLM with basic information about their persona and the company.
@@
The model can eventually decide to (1) deny, (2) approve half of the amount, or (3) provide the full amount. We consider both the decisions (2) and (3) as misconduct.
@@ Appendix B Pressure variables
Appendix C Models
C.1 Models employed
C.1.1 Closed access models
The snapshots of the OpenAI models used in the experiments are:
  • gpt-4o-mini-2024-07-18
  • gpt-4o-2024-05-13
  • o1-preview-2024-09-12
  • o1-mini-2024-09-12
  • o3-mini-2025-01-31
  • gpt-4-turbo-2024-04-09
  • gpt-3.5-turbo-0125
For Claude 3 Haiku, the snapshot used is claude-3-haiku-20240307; the claude-3-5-sonnet-20240620 snapshot has been used for Sonnet 3.5; while for Claude 3.5 Haiku we use claude-3-5-haiku-20241022.
C.1.2 Open access models
Our model selection contains two open-access models: phi-3.5-mini (Abdin et al., 2024) and llama-3.1-8b (Dubey et al., 2024). The model weights were accessed through the official Huggingface repositories. We use the instruct version of both models, and format the prompts with the provided chat templates to ensure correct text generation.
Appendix D Choice of sample size
$\textrm{SE}_{\hat{p}}^{\text{max}}(N)=\max_{p}\sqrt{p(1-p)/N}.$ (2)
Appendix E Additional results
E.1 Table of parameters
In Table 4 and 5 we report the results of the logistic regression analysis for all LLMs considered. The two tables respectively indicate the parameters of the model and the corresponding odds ratios. Parameters can be positive or negative, a positive (negative) value indicates that a given parameter value decreases (increases) the probability of misalignment. On the other hand, odds ratios are always positive and represent the ratios of the misalignment probabilities with and without the use of a specific prompt variable. The short names in the ‘variable’ column indicate the type of pressure exerted (e.g., ‘risk’), and whether the expected sign of the coefficient is positive (e.g., ‘risk+’) or negative (e.g., ‘risk-’).
E.2 Results with T=0.1
Relationships with sycophancy benchmarks.
Sycophancy is an undesirable behavior exhibited by models when they align their responses and opinions with the user’s perspective, regardless of its correctness (Perez et al., 2023). Sharma et al. (2023) suggests that this tendency may be more marked in LLMs that have been trained to follow human feedback. In order to compare the occurrence of this behavior to the misalignment rate found in our experiment, we measure sycophancy using the lm-exp-sycophancy (Rimsky, 2023) and Opinion Pairs (Huang et al., 2024) datasets. As shown in Figure 8, we do not find any statistically significant relationship with our misalignment metric.
Appendix F Robustness checks on the logistic regression results
Ordinal logistic regression.
Autoregressive logistic regression.
We hypothesize that the autoregressive nature of LLMs implies that, generally speaking, dependencies may exist among the variables, even with respect to the order in which they are presented in the prompt. To strengthen our results, we repeated the regression exercise using an autoregressive extension of logistic regression and confirmed that the qualitative outcomes were equivalent to the original results. Specifically, we used a recurrent neural network (RNN) implementing the following operations. First, the input variables are passed through a fully connected layer with a one-dimensional output. Then, this one-dimensional output is summed to the one-dimensional hidden space (a kind of “misalignment state”) and passed to a tanh activation function to generate a new hidden space. Finally, the misalignment state is multiplied by a parameter and passed through a sigmoid function to predict the misalignment probability. An illustration of this architecture is provided in Figure 9. We train the network’s parameters using a cross-entropy loss between the misalignment decision made by the LLM and the final predicted misalignment probability $p_{7}$. We train for each model for 20 epochs using a batch size of 32, an Adam optimizer and a weight decay of $10^{-4}$. This model, which we can consider a kind of “autoregressive logistic regression”, yields results that are qualitatively equivalent to those presented in the main text, as shown in Figure 10 and in Table 8. The RNNs model the probability of misalignment as a function of the prompt variable and the previously computed hidden misalignment state. The marginal effect that each prompt variable has on the probability of misalignment is depicted in Figure 11 for a subset of models. The figure illustrates the different baseline propensities to misalign across models, as well as the asymmetric effect that each prompt variable can have on $p$.
[…]
Appendix G Analysis of LLM Prompts
In this section, we analyze how the models respond to our simulated scenario by identifying the used terms and categorizing them. We define five categories: ’misappropriation’, ’legal’, ’illegal’, ’ethical’, and ’unethical’. For each category, we provide a list of several related terms in Table 2. It is important to note that these five categories are not explicitly mentioned in our input prompts.
Category Terms
@@ Appendix G Analysis of LLM Prompts
variable gpt-3.5-turbo gpt-4-turbo claude-3-haiku claude-3.5-haiku claude-son-3.5 gpt-4o gpt-4o-mini llama3.1-8b phi3.5-mini o1-mini o1-preview o3-mini
risk+ 0.14∗∗∗ 1.71∗∗∗ 0.26∗∗∗ 3.12∗∗∗ 5.20 ∗∗∗ 1.99∗∗∗ 1.22∗∗∗ 0.90∗∗∗ 0.34∗∗∗ 0.88∗∗∗ 1.54∗∗∗ 1.09∗∗∗
(0.03) (0.03) (0.02) (0.05) (0.06) (0.04) (0.03) (0.04) (0.03) (0.04) (0.04) (0.03)
risk- -0.12∗∗∗ -0.43∗∗∗ -0.23∗∗∗ -0.18∗∗∗ -2.42 ∗∗∗ -1.05∗∗∗ -0.97∗∗∗ -0.18∗∗∗ -0.31∗∗∗ -0.72∗∗∗ -0.77∗∗∗ -0.85∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.03) (0.02) (0.03) (0.05) (0.03)
reg+ -0.13∗∗∗ 0.05 -0.12∗∗∗ 0.22∗∗∗ 0.34∗∗∗ 0.05 -0.05 0.12∗∗∗ 0.05∗∗ 0.01 0.89 ∗∗∗ 0.01
(0.03) (0.03) (0.02) (0.04) (0.03) (0.04) (0.03) (0.04) (0.03) (0.03) (0.03) (0.03)
reg- -0.36∗∗∗ -1.80∗∗∗ -0.49∗∗∗ -2.27∗∗∗ -3.82 ∗∗∗ -1.72∗∗∗ -0.39∗∗∗ -0.41∗∗∗ -0.68∗∗∗ -0.70∗∗∗ -2.34∗∗∗ -1.01∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.05) (0.03) (0.03) (0.04) (0.02) (0.03) (0.06) (0.03)
loan+ -0.01 0.38 ∗∗∗ 0.11∗∗∗ 0.15∗∗∗ 0.15∗∗∗ 0.27∗∗∗ 0.16∗∗∗ 0.07 0.07∗∗∗ -0.05 0.22∗∗∗ 0.29∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03)
loan- -0.32∗∗∗ -0.27∗∗∗ -0.32∗∗∗ -0.42∗∗∗ -0.27∗∗∗ -0.66∗∗∗ -0.36∗∗∗ -0.13∗∗∗ -0.30∗∗∗ -0.21∗∗∗ -0.26∗∗∗ -0.99 ∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.04) (0.02) (0.03) (0.04) (0.03)
gov+ -0.23∗∗∗ -0.17∗∗∗ -0.58∗∗∗ -0.78∗∗∗ -0.44∗∗∗ -0.32∗∗∗ -0.31∗∗∗ -0.25∗∗∗ -0.27∗∗∗ -0.09∗∗∗ -0.08 ∗∗ -0.18∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03)
gov- 0.02 0.17∗∗∗ 0.10∗∗∗ 0.30∗∗∗ 0.27∗∗∗ -0.15∗∗∗ 0.08∗∗∗ -0.00 -0.09∗∗∗ 0.16∗∗∗ -0.45 ∗∗∗ 0.51∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03)
trust+ 0.41∗∗∗ 1.38∗∗∗ -0.09∗∗∗ 2.13 ∗∗∗ 1.44∗∗∗ 1.25∗∗∗ 0.86∗∗∗ 0.72∗∗∗ 0.20∗∗∗ 0.35∗∗∗ 0.67∗∗∗ 1.54∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.04) (0.03) (0.05) (0.03) (0.03) (0.03) (0.04)
trust- -0.51∗∗∗ -0.59∗∗∗ -0.66∗∗∗ -0.46∗∗∗ -0.80∗∗∗ -0.81∗∗∗ -0.92∗∗∗ -0.78∗∗∗ -0.45∗∗∗ -0.52∗∗∗ -0.48∗∗∗ -1.08 ∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.03) (0.02) (0.03) (0.04) (0.03)
outlook+ 0.07∗∗ 0.11∗∗∗ 0.08∗∗∗ 0.31 ∗∗∗ -0.01 -0.18∗∗∗ 0.14∗∗∗ 0.15∗∗∗ 0.10∗∗∗ 0.04 -0.15∗∗∗ 0.08∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.04) (0.03)
outlook- 0.22∗∗∗ 0.08∗∗∗ -0.02 0.51∗∗∗ 0.18∗∗∗ 0.04 0.19∗∗∗ 0.04 -0.04 0.10∗∗∗ -0.21 ∗∗∗ 0.16∗∗∗
(0.03) (0.03) (0.02) (0.03) (0.04) (0.03) (0.03) (0.04) (0.02) (0.03) (0.04) (0.03)
profitexp+ 1.22∗∗∗ 1.84∗∗∗ 0.99∗∗∗ 1.50∗∗∗ 0.97∗∗∗ 1.02∗∗∗ 1.48∗∗∗ 1.01∗∗∗ 1.01∗∗∗ 0.90∗∗∗ 0.49∗∗∗ 2.47 ∗∗∗
(0.03) (0.03) (0.02) (0.04) (0.04) (0.04) (0.03) (0.04) (0.03) (0.04) (0.03) (0.05)
profitexp- 0.05∗∗ -3.40∗∗∗ -0.62∗∗∗ -1.31∗∗∗ -3.42 ∗∗∗ -2.50∗∗∗ -1.27∗∗∗ 0.01 -0.60∗∗∗ -1.59∗∗∗ -0.67∗∗∗ -1.32∗∗∗
(0.03) (0.04) (0.02) (0.03) (0.05) (0.03) (0.03) (0.03) (0.02) (0.03) (0.04) (0.03)
constant 1.38∗∗∗ -0.51∗∗∗ 0.77∗∗∗ 1.62∗∗∗ 0.47∗∗∗ 3.20 ∗∗∗ -0.40∗∗∗ 1.95∗∗∗ 1.41∗∗∗ 2.67∗∗∗ -2.38∗∗∗ 1.88∗∗∗
(0.04) (0.05) (0.04) (0.05) (0.06) (0.06) (0.04) (0.06) (0.04) (0.05) (0.06) (0.05)
$N$ 52130 54356 54447 54668 52852 54537 54574 46273 53584 54367 54301 54675
R2 0.07 0.45 0.11 0.41 0.63 0.40 0.28 0.10 0.10 0.20 0.27 0.36
@@
variable gpt-3.5-turbo gpt-4-turbo claude-3-haiku claude-3.5-haiku claude-son-3.5 gpt-4o gpt-4o-mini llama3.1-8b phi3.5-mini o1-mini o1-preview o3-mini
risk+ 1.15∗∗∗ 5.55∗∗∗ 1.30∗∗∗ 22.57∗∗∗ 181.16 ∗∗∗ 7.28∗∗∗ 3.37∗∗∗ 2.46∗∗∗ 1.40∗∗∗ 2.40∗∗∗ 4.64∗∗∗ 2.97∗∗∗
(0.03) (0.18) (0.03) (1.11) (10.46) (0.30) (0.09) (0.10) (0.04) (0.09) (0.16) (0.10)
risk- 0.89∗∗∗ 0.65∗∗∗ 0.80∗∗∗ 0.83∗∗∗ 0.09 ∗∗∗ 0.35∗∗∗ 0.38∗∗∗ 0.83∗∗∗ 0.73∗∗∗ 0.49∗∗∗ 0.46∗∗∗ 0.43∗∗∗
(0.02) (0.02) (0.02) (0.02) (0.00) (0.01) (0.01) (0.03) (0.02) (0.01) (0.02) (0.01)
reg+ 0.88∗∗∗ 1.05 0.88∗∗∗ 1.25∗∗∗ 1.41∗∗∗ 1.05 0.95 1.13∗∗∗ 1.05∗∗ 1.01 2.44 ∗∗∗ 1.01
(0.02) (0.03) (0.02) (0.05) (0.05) (0.04) (0.02) (0.04) (0.03) (0.03) (0.08) (0.03)
reg- 0.70∗∗∗ 0.16∗∗∗ 0.62∗∗∗ 0.10∗∗∗ 0.02 ∗∗∗ 0.18∗∗∗ 0.68∗∗∗ 0.66∗∗∗ 0.51∗∗∗ 0.50∗∗∗ 0.10∗∗∗ 0.36∗∗∗
(0.02) (0.01) (0.01) (0.00) (0.00) (0.01) (0.02) (0.02) (0.01) (0.02) (0.01) (0.01)
loan+ 0.99 1.46 ∗∗∗ 1.12∗∗∗ 1.17∗∗∗ 1.16∗∗∗ 1.31∗∗∗ 1.17∗∗∗ 1.07 1.07∗∗∗ 0.95 1.24∗∗∗ 1.33∗∗∗
(0.03) (0.04) (0.03) (0.04) (0.04) (0.05) (0.03) (0.04) (0.03) (0.03) (0.04) (0.04)
loan- 0.72∗∗∗ 0.77∗∗∗ 0.73∗∗∗ 0.66∗∗∗ 0.76∗∗∗ 0.52∗∗∗ 0.69∗∗∗ 0.88∗∗∗ 0.74∗∗∗ 0.81∗∗∗ 0.77∗∗∗ 0.37 ∗∗∗
(0.02) (0.02) (0.02) (0.02) (0.03) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03) (0.01)
gov+ 0.80∗∗∗ 0.85∗∗∗ 0.56∗∗∗ 0.46∗∗∗ 0.65∗∗∗ 0.73∗∗∗ 0.73∗∗∗ 0.78∗∗∗ 0.76∗∗∗ 0.91∗∗∗ 0.93 ∗∗ 0.84∗∗∗
(0.02) (0.03) (0.01) (0.02) (0.02) (0.02) (0.02) (0.03) (0.02) (0.03) (0.03) (0.03)
gov- 1.02 1.19∗∗∗ 1.10∗∗∗ 1.35∗∗∗ 1.31∗∗∗ 0.86∗∗∗ 1.08∗∗∗ 1.00 0.91∗∗∗ 1.17∗∗∗ 0.64 ∗∗∗ 1.67∗∗∗
(0.03) (0.04) (0.03) (0.05) (0.05) (0.03) (0.03) (0.04) (0.02) (0.04) (0.02) (0.05)
trust+ 1.51∗∗∗ 3.96∗∗∗ 0.91∗∗∗ 8.39 ∗∗∗ 4.23∗∗∗ 3.51∗∗∗ 2.36∗∗∗ 2.05∗∗∗ 1.22∗∗∗ 1.41∗∗∗ 1.96∗∗∗ 4.69∗∗∗
(0.05) (0.13) (0.02) (0.35) (0.17) (0.13) (0.06) (0.09) (0.03) (0.05) (0.07) (0.17)
trust- 0.60∗∗∗ 0.55∗∗∗ 0.52∗∗∗ 0.63∗∗∗ 0.45∗∗∗ 0.44∗∗∗ 0.40∗∗∗ 0.46∗∗∗ 0.64∗∗∗ 0.60∗∗∗ 0.62∗∗∗ 0.34 ∗∗∗
(0.02) (0.02) (0.01) (0.02) (0.02) (0.01) (0.01) (0.02) (0.02) (0.02) (0.02) (0.01)
outlook+ 1.07∗∗ 1.11∗∗∗ 1.08∗∗∗ 1.36 ∗∗∗ 0.99 0.83∗∗∗ 1.15∗∗∗ 1.16∗∗∗ 1.11∗∗∗ 1.04 0.86∗∗∗ 1.08∗∗
(0.03) (0.03) (0.02) (0.05) (0.04) (0.03) (0.03) (0.04) (0.03) (0.03) (0.03) (0.03)
outlook- 1.25∗∗∗ 1.09∗∗∗ 0.99 1.66∗∗∗ 1.20∗∗∗ 1.04 1.21∗∗∗ 1.04 0.96 1.11∗∗∗ 0.81 ∗∗∗ 1.17∗∗∗
(0.03) (0.03) (0.02) (0.06) (0.04) (0.04) (0.03) (0.04) (0.02) (0.04) (0.03) (0.04)
profitexp+ 3.39∗∗∗ 6.33∗∗∗ 2.70∗∗∗ 4.46∗∗∗ 2.65∗∗∗ 2.79∗∗∗ 4.37∗∗∗ 2.74∗∗∗ 2.75∗∗∗ 2.46∗∗∗ 1.63∗∗∗ 11.84 ∗∗∗
(0.11) (0.18) (0.06) (0.18) (0.10) (0.12) (0.11) (0.12) (0.08) (0.11) (0.06) (0.55)
profitexp- 1.05∗∗ 0.03∗∗∗ 0.54∗∗∗ 0.27∗∗∗ 0.03 ∗∗∗ 0.08∗∗∗ 0.28∗∗∗ 1.01 0.55∗∗∗ 0.20∗∗∗ 0.51∗∗∗ 0.27∗∗∗
(0.03) (0.00) (0.01) (0.01) (0.00) (0.00) (0.01) (0.03) (0.01) (0.01) (0.02) (0.01)
constant 3.99∗∗∗ 0.60∗∗∗ 2.16∗∗∗ 5.04∗∗∗ 1.60∗∗∗ 24.50 ∗∗∗ 0.67∗∗∗ 7.02∗∗∗ 4.10∗∗∗ 14.44∗∗∗ 0.09∗∗∗ 6.54∗∗∗
(0.17) (0.03) (0.08) (0.26) (0.09) (1.40) (0.03) (0.41) (0.16) (0.77) (0.01) (0.32)
$N$ 52130 54356 54447 54668 52852 54537 54574 46273 53584 54367 54301 54675
R2 0.07 0.45 0.11 0.41 0.63 0.40 0.28 0.10 0.10 0.20 0.27 0.36